# Alpha-Particle Monopole Form Factors with Ab Initio No-Core Shell Model

### v1

@@ III Results and discussion
We start with calculations of the elastic 4He form factor. In Fig. 1 we present the point-proton density distribution of the 4He ground state for with the same MeV. We observe good exponential tail for the proton density and its robust convergence with respect to . The results for are nearly indistinguishable in Fig. 1. We performed similar calculations for MeV and found also robust convergence. For simplicity we show in Fig. 1 only the results of MeV at . We find that the results of and MeV at are in good agreement with each other.
Based on the point-proton density distribution shown in Fig. 1, we can obtain the elastic form factor straightforwardly as we mention in Sec. II. In Fig. 2 we present the elastic form factor of the 4He ground state for with the same MeV and for with MeV. As we explain in Eq. (3), the elastic form factor is unity at fm-2. The NCSM results show good convergence with respect to . The converged results are in good agreement with the experimental data, which is not surprising, since the Daejeon16 interaction gives the 4He point-proton radius of fm, which is very close to the experimental value of 1.484(5) fm. Our results are also in good agreement with the EIHH calculations with the Argonne V18 (AV18) potential plus the Urbana IX (UIX) 3 force and Chiral interactions ( interaction at N3LO and 3 interaction at N2LO) in Ref. [9], which give the 4He point-proton radius of and fm, respectively. Neglecting two-body currents and further relativistic corrections, the experimental value of the point-proton radius is obtained using the experimental charge radius fm [42] via the following expression [43]:

### v2

@@ III Results and discussion
We start with calculations of the elastic 4He form factor. In Fig. 1 (a) we present the point-proton density distribution of the 4He ground state for with the same MeV. We observe good exponential tail for the proton density and its robust convergence with respect to . The results for are nearly indistinguishable in Fig. 1 (a). We performed similar calculations for MeV and found also robust convergence. For simplicity we show in Fig. 1 (a) only the results of MeV at . We find that the results of and MeV at are in good agreement with each other.
Based on the point-proton density distribution shown in Fig. 1 (a), we can obtain the elastic form factor straightforwardly as we mention in Sec. II. In Fig. 1 (b) we present the elastic form factor of the 4He ground state for with the same MeV and for with MeV. As we explain in Eq. (3), the elastic form factor is unity at fm-2. The NCSM results show good convergence with respect to . The converged results are in good agreement with the experimental data, which is not surprising, since the Daejeon16 interaction gives the 4He point-proton radius of fm, which is very close to the experimental value of 1.484(5) fm. Our results are also in good agreement with the EIHH calculations with the Argonne V18 (AV18) potential plus the Urbana IX (UIX) 3 force and Chiral interactions ( interaction at N3LO and 3 interaction at N2LO) in Ref. [9], which give the 4He point-proton radius of and fm, respectively. Neglecting two-body currents and further relativistic corrections, the experimental value of the point-proton radius is obtained using the experimental charge radius fm [42] via the following expression [43]:
@@
We noticed that the point-proton density and the elastic form factor have been calculated with the NCSM using the Daejeon16 interaction at in Ref. [45]. The proton form factor adopted in Ref. [45], which corresponds to the proton charge radius of fm, was taken from Ref. [46]. Our results are consistent with those in Ref. [45]. We illustrate the convergence of the these two quantities as the basis size increases.
The robust convergence and good description of the experimental data demonstrated in Fig. 1 support the reliability of our calculations of the monopole transition form factor of 4He.